### v1

@@ Abstract.
We study the class of ${\mathcal{K}}$-Lorentzian polynomials, a generalization of the distinguished class of Lorentzian polynomials. As shown in [BD24], the set of ${\mathcal{K}}$-Lorentzian polynomials is equivalent to the set of ${\mathcal{K}}$-completely log-concave (aka ${\mathcal{K}}$-CLC) forms. Throughout this paper, we interchangeably use the terms ${\mathcal{K}}$-Lorentzian polynomials for the homogeneous setting and ${\mathcal{K}}$-CLC polynomials for the non-homogeneous setting. By introducing an alternative definition of ${\mathcal{K}}$-CLC polynomials through univariate restrictions, we establish that any strictly ${\mathcal{K}}$-CLC polynomial of degree $d\leq 4$ is Hurwitz-stable polynomial over ${\mathcal{K}}$. Additionally, we characterize the conditions under which a strictly ${\mathcal{K}}$-CLC of degree $d\geq 5$ is Hurwitz-stable over ${\mathcal{K}}$. Furthermore, we associate the largest possible proper cone, denoted by ${\mathcal{K}}(f,v)$, with a given ${\mathcal{K}}$-Lorentzian polynomial $f$ in the direction $v\in\operatorname{int}{\mathcal{K}}$. Finally, we investigate applications of ${\mathcal{K}}$-CLC polynomials in the stability analysis of evolution variational inequalities (EVI) dynamical systems governed by differential equations and inequality constraints.
@@
Lorentzian polynomials, convex cone, Hurwitz-stable, Rayleigh Difference polynomial, log-concavity, semipositive cone, EVI dynamical system
@@ 1. Introduction
${\mathcal{K}}$-Lorentzian polynomials are introduced and studied by Brändén and Leake in [BL23a], and by Blekherman and Dey in [BD24]. Our approach in [BD24] not only generalizes Lorentzian polynomials introduced by June Huh and Petter Brändén in [BH20] and simultaneously by Anari, Oveis Gharan and Vinzant [AGV21] (under the name of completely log-concave polynomials), but also provides a unified framework for extending this class. In [BD24], we established a link between ${\mathcal{K}}$-Lorentzian polynomials and ${\mathcal{K}}$-nonnegative matrices, which arise in the context of the generalized Perron-Frobenius Theorem. Leveraging this connection, we presented a further characterization of ${\mathcal{K}}$-Lorentzian polynomials over self-dual cones, yielding an additional necessary and sufficient condition for Lorentzian polynomials over the nonnegative orthant.
We observe that for any ${\mathcal{K}}$-Lorentzian polynomial over self-dual cone ${\mathcal{K}}$, the nonsingular Hessian matrices evaluated at points $a\in{\mathcal{K}}$ satisfy the generalized Perron-Frobenius Theorem. We introduce an alternative definition of ${\mathcal{K}}$-CLC polynomials by restricting them to lines, which enables the analysis of log-concavity in the coefficients of the resulting univariate polynomials. Furthermore, leveraging the structure of the coefficients, we show that any strictly ${\mathcal{K}}$-CLC polynomial of degree $d\leq 4$ is Hurwitz-stable over ${\mathcal{K}}$, [see Theorem 4.20]. Moreover, for degree $d\geq 5$, we characterize the class of strictly ${\mathcal{K}}$-CLC polynomials which are Hurwitz-stable over ${\mathcal{K}}$ using a known result to characterize Hurwitz-stability for univariate polynomials of degree $d\geq 5$, [see Theorem 4.21]. An example of a strictly CLC polynomial of $d=5$ that is not Hurwitz stable is provided. This plays a crucial role in stability analysis of a dynamical system when trajectories are constrained on certain regions.
Throughout this article, unless otherwise stated, we assume that ${\mathcal{K}}$ is a proper convex cone. A natural progression is to determine an appropriate ${\mathcal{K}}$ for a given ${\mathcal{K}}$-Lorentzian polynomial, similar to the way hyperbolicity cones correspond to hyperbolic polynomials. To achieve this, we identify the largest possible proper cone w.r.t a direction $v$, denoted by ${\mathcal{K}}(f,v)$, associated with a ${\mathcal{K}}$-Lorentzian candidate. We provide a characterization of the interior of the closed cone ${\mathcal{K}}(f,v)$, [see Theorem 5.5] and demonstrate that if $f$ is ${\mathcal{K}}$-Lorentzian on ${\mathcal{K}}(f,v)$, then ${\mathcal{K}}(f,v)$ is a convex set, [see Theorem 5.7]. Notably, when $f$ is a hyperbolic polynomial, ${\mathcal{K}}(f,v)$ corresponds to the hyperbolicity cone of $f$ in the direction $v$. The Rayleigh -Difference polynomial, defined as $\Delta_{i,j}f(x):=\partial_{i}f\partial_{j}f-f\partial_{i}\partial_{j}f$, plays a significant role in optimization and sampling algorithms (fast algorithm in MCMC methods), especially in high-dimensional, non-convex settings. In statistics, many distributions (e.g., Gaussian, exponential) are log-concave. The nonnegative condition of the Rayleigh difference polynomial ensures the preservation of log-concavity under marginalization or convolution. We demonstrate that for a ${\mathcal{K}}$-Lorentzian polynomial $f$, the Rayleigh difference polynomials $D_{w}f.D_{v}f-f\cdot D_{w}D_{v}f$ are nonnegative over ${\mathcal{K}}$ for all $v,w\in{\mathcal{K}}$, [see Theorem 5.14]. An additional area of exploration involves the semipositive cone, obtained as the intersection of a hyperbolicity cone with the nonnegative orthant, for a hyperbolic generating polynomial $f_{A}$ associated with a nonsingular semipositive matrix $A$, [see Proposition 5.15]. This class of polynomials facilitates the characterization of the intersection cone, as the ${\mathcal{K}}$-semipositive cone, [see Theorem 5.21].
The Hurwitz-stability of strictly CLC polynomial $f(t)$ over the interval $t\geq 0$ establishes a connection to stable LTI dynamical systems of the form $\frac{dx}{dt}=Ax$, where the characteristic polynomial $\chi(A)=f(t)$, [see Theorem 6.2] and Theorem 6.4]. Moreover, we introduce a method to assess the (Lyapunov-type) stability w.r.t ${\mathcal{K}}$ of certain dynamical systems through ${\mathcal{K}}$-Lorentzian polynomials. This class is known as the Evolution Variational Inequality (EVI), and characterized by a combination of differential equations and inequality constraints. These systems describe the evolution of a state $x(t)$ in time, constrained by a variational inequality associated with a convex set and a functional that defines the system dynamics. Specifically, for EVI systems, we show that even if the system lacks global (asymptotic) stability, it can still exhibit (asymptotic) stability w.r.t ${\mathcal{K}}$ via ${\mathcal{K}}$-CLC polynomials, [see Theorem 6.18 and Theorem 6.23]. Additionally, ${\mathcal{K}}$-Lorentzian polynomials can be employed to identify regions where stability is achievable. EVIs often arise as the continuous-time counterpart of optimization problems constrained by convex sets. For example, they can model gradient flows constrained by feasibility conditions, where the goal is to minimize a convex function $f(x)$ subject to $x\in{\mathcal{K}}$. EVIs are extensively used in applied mathematics and various scientific disciplines, with applications spanning control theory (e.g., when state variables or control inputs are constrained), game theory (such as saddle-point dynamics and Nash equilibrium computation), machine learning (e.g., gradient descent flows with constraints), and traffic and network flow analysis. For additional applications, see [GB04] and the references therein.
The rest of the paper is structured as follows: In Section 2 we recall the definitions and essential theorems that will be utilized in the following sections. In Section 3, we develop several key properties of ${\mathcal{K}}$-Lorentzian polynomials. Notably, we introduce an alternative definition of ${\mathcal{K}}$-CLC polynomials (which need not be forms) by restricting them to specific lines, allowing us to derive important necessary conditions for ${\mathcal{K}}$-CLC polynomials. One such condition highlights the log-concavity properties of the coefficients of the univariate restricted polynomials. Section 4 explores the connection between Hurwitz-stability over ${\mathcal{K}}$ and strictly ${\mathcal{K}}$-CLC polynomials, noting that not all strictly ${\mathcal{K}}$-CLC polynomials of degree $d\geq 5$ are Hurwitz-stable, and examining cases where they are. In Section 5, the largest proper cone ${\mathcal{K}}(f,v)$ with a given ${\mathcal{K}}$-CLC polynomial $f$ is introduced, along with nonnegativity results for Rayleigh difference polynomials and discussions on ${\mathcal{K}}$-semipositive cones. Finally, Section 6 demonstrates how stability analysis of EVI systems is achieved through ${\mathcal{K}}$-CLC polynomials.
@@ 2. State of the Art
Let ${\mathbb{R}}[x]$ represent the space of $n$-variate polynomials over ${\mathbb{R}}$, and ${\mathbb{R}}[x]_{\leq d}$ represent the space of $n$-variate polynomials over ${\mathbb{R}}$ with degree at most $d$ and ${\mathbb{R}}[x]^{d}_{n}$ denote the space of real homogeneous polynomials (aka forms) in $n$ variables of degree $d$. A polynomial $f\in{\mathbb{R}}[x]$ is said to be (strictly) log-concave at a point $a\in{\mathbb{R}}^{n}$ if $f(a)>0$, and $\log f$ is a (strictly) concave function at $a$, i.e. the Hessian of $\log f$ is negative semidefinite (negative definite) at $a$. A polynomial $f\in{\mathbb{R}}[x]$ is log-concave on a proper cone ${\mathcal{K}}\subset{\mathbb{R}}^{n}$ if $f$ is log-concave at every point of the interior of ${\mathcal{K}}$. By convention, the zero polynomial is log-concave (but not strictly log-concave) at all points of ${\mathbb{R}}^{n}$.
If ${\mathcal{K}}={\mathcal{K}}^{\ast}$, a self-dual cone, then $K$ is closed, pointed, and full dimensional[BF76]. We denote quadratic forms by a lower case letter, and the matrix of the quadratic form by the corresponding upper case letter, i.e. if $q$ is a quadratic form, then its matrix is $Q$ and $q(x)=x^{t}Qx$.
@@
[BD24] A polynomial (form) $f\in{\mathbb{R}}[x]_{\leq d}$ is called a ${\mathcal{K}}$-completely log-concave aka ${\mathcal{K}}$-CLC (form) on a proper convex cone ${\mathcal{K}}$ if for any choice of $a_{1},\dots,a_{m}\in{\mathcal{K}}$, with $m\leq d$, we have that $D_{a_{1}}\dots D_{a_{m}}f$ is log-concave on $\operatorname{int}{\mathcal{K}}$. A polynomial (form) $f\in{\mathbb{R}}[x]$ is strictly ${\mathcal{K}}$-CLC if for any choice of $a_{1},\dots,a_{m}\in{\mathcal{K}}$, with $m\leq d$, $D_{a_{1}}\dots D_{a_{m}}f$ is strictly log-concave on all points of ${\mathcal{K}}$.
@@
[BD24] Let ${\mathcal{K}}$ be a proper convex cone. A form $f\in{\mathbb{R}}[x]_{n}^{d}$ of degree $d\geq 2$ is said to be ${\mathcal{K}}$-Lorentzian if for any $a_{1},\dots,a_{d-2}\in\operatorname{int}{\mathcal{K}}$, the quadratic form $q=D_{a_{1}}\dots D_{a_{d-2}}f$ satisfies the following conditions:
@@
Let ${\mathcal{K}}$ be a proper convex cone in ${\mathbb{R}}^{n}$. An $n\times n$ matrix $A$ is ${\mathcal{K}}$-nonnegative if $Ax\in{\mathcal{K}}$ for any $x\in{\mathcal{K}}$, i.e. $A({\mathcal{K}})\subseteq{\mathcal{K}}$. A matrix $A$ is ${\mathcal{K}}$-positive if $Ax\in\operatorname{int}{\mathcal{K}}$ for any nonzero $x\in{\mathcal{K}}$. A ${\mathcal{K}}$-nonnegative matrix $A$ is called ${\mathcal{K}}$-irreducible if $A$ leaves no (proper) face of ${\mathcal{K}}$ invariant. A ${\mathcal{K}}$-nonnegative matrix which is not ${\mathcal{K}}$-irreducible is called ${\mathcal{K}}$-reducible, cf. [Gan98], [Van68].
@@
[BD24, Lemma 3.1] A quadratic form $q(x)=x^{t}Qx\in{\mathbb{R}}[x]_{n}^{2}$ is ${\mathcal{K}}$-Lorentzian if and only if the matrix $Q$ of $q$ has exactly one positive eigenvalue and one of the following conditions holds:
@@
Similar to copositive matrices over the nonnegative orthant, we define ${\mathcal{K}}$-copositive matrices over ${\mathcal{K}}$, see [GB04]. A matrix $Q$ is (strictly) ${\mathcal{K}}$-copositive if $x^{t}Qx(>0)\geq 0$ for all $x\in{\mathcal{K}}$.
@@
[BD24, Cor 3.8] Let ${\mathcal{K}}$ be a self-dual cone. Then the quadratic form $q=x^{t}Qx$ is ${\mathcal{K}}$-Lorentzian if and only if the matrix $Q$ of $q$ has exactly one positive eigenvalue and $Q$ is ${\mathcal{K}}$-nonnegative.
@@
[BD24, Cor 3:10] Let ${\mathcal{K}}\subset{\mathbb{R}}^{n}$ be a self-dual cone. Then $q=x^{t}Qx$ is ${\mathcal{K}}$-Lorentzian if and only if the matrix $Q$ of $q$ has exactly one positive eigenvalue and $Q$ is either nonsingular and ${\mathcal{K}}$-irreducible, or singular and ${\mathcal{K}}$-nonnegative.

### v2

@@ Abstract.
We investigate the class of $\mathcal{K}$}-Lorentzian polynomials, which generalizes the well-known class of Lorentzian polynomials. As demonstrated in [GPlorentzian], these polynomials are equivalent to $\mathcal{K}$}-completely log-concave ($\mathcal{K}$}-CLC) forms. Exploiting the structure of directional derivatives and Hessians, we prove that the Hessian matrices of $\mathcal{K}$}-CLC forms satisfy a generalized Perron–Frobenius property over a self-dual cone $\mathcal{K}$}: they admit a distinguished eigenvector in $\operatorname{int}{\mathcal{K}}$. We introduce a line-restriction perspective on $\mathcal{K}$}-CLC. This leads to a hierarchy of log-concavity and Newton-type log-concavity inequalities for the coefficients of univariate restrictions. This provides useful necessary conditions for $\mathcal{K}$}-CLC in terms of directional derivatives.
As a key application, we derive degree-dependent criteria for Hurwitz stability of $\mathcal{K}$}-CLC univariate restrictions. In particular, we prove that all strictly $\mathcal{K}$}-Lorentzian polynomials of degree $d\leq 4$ are Hurwitz-stable over $\mathcal{K}$}, and we obtain refined inequalities that give sufficient conditions for Hurwitz stability in degrees $d\geq 5$. These results recover and extend known stability criteria in the classical Lorentzian case ${\mathcal{K}}={\mathbb{R}}^{n}_{+}$ to the general conic setting. Moreover, these Hurwitz–stability and Newton-type log-concavity coefficient criteria yield simple sufficient conditions for asymptotic stability of linear time-invariant systems via their characteristic polynomials.
@@
$\mathcal{K}$}-Lorentzian polynomials, $\mathcal{K}$}-completely log-concave polynomials, convex cones, log-concavity, Newton-type inequalities, Perron–Frobenius theorem, Hurwitz stability, linear time-invariant (LTI) systems
@@ 1. Introduction
$\mathcal{K}$}-Lorentzian polynomials were introduced and studied by Brändén and Leake in [Leake], and by Blekherman and Dey in [GPlorentzian]. The approach in [GPlorentzian] not only generalizes Lorentzian polynomials introduced by Huh and Brändén [Brandenlorentzian] and, independently, by Anari, Oveis Gharan and Vinzant under the name of completely log-concave polynomials [Cynthialog], but also provides a unified framework for extending this class to arbitrary proper convex cones.
In this work we use the term $\mathcal{K}$}-Lorentzian for the homogeneous setting and $\mathcal{K}$}-completely log-concave ($\mathcal{K}$}-CLC) for the possibly non-homogeneous setting. A central result of [GPlorentzian] is that, for any proper cone $\mathcal{K}$}, the class of $\mathcal{K}$}-Lorentzian forms coincides with the class of $\mathcal{K}$}-CLC forms of the same degree; see Theorem˜2.9. This gives a powerful bridge between convex-analytic log-concavity on cones and the algebraic signature conditions that define Lorentzian behavior via iterated directional derivatives and quadratic forms.
Our first goal in this part is to refine the structural picture of $\mathcal{K}$}-Lorentzian and $\mathcal{K}$}-CLC polynomials. For homogeneous forms, we show that the Hessian matrices evaluated on $\mathcal{K}$} have a very rigid eigenvalue structure: they have exactly one positive eigenvalue, and, over self-dual cones, satisfy a generalized Perron–Frobenius property (Theorem˜3.6, Theorem˜3.7, Corollary˜3.8). This connects $\mathcal{K}$}-Lorentzian polynomials to $\mathcal{K}$}-nonnegative and $\mathcal{K}$}-irreducible matrices, and to the cone-theoretic Perron–Frobenius theory developed in [Kreinlinear, Barkeralgebraic].
Next, we introduce a linewise notion of $\mathcal{K}$}-complete log-concavity: for each $x\in\operatorname{int}{\mathcal{K}}$ and $v\in{\mathcal{K}}$, we consider the univariate restriction
$t\longmapsto f(x+tv),$
and require all iterated derivatives in $t$ to be log-concave on the appropriate interval. We show that $\mathcal{K}$}-CLC in the sense of Definition˜2.1 implies linewise $\mathcal{K}$}-CLC (Proposition˜3.13). Combined with sharp univariate characterizations of complete log-concavity, this yields log-concavity of the factorial-weighted coefficient sequences of $f(x+tv)$ and, consequently, log-concavity of the directional derivative sequence
$\bigl\{f(x),D_{v}f(x),D_{v}^{2}f(x),\dots,D_{v}^{d}f(x)\bigr\}$
for any $x,v\in\operatorname{int}{\mathcal{K}}$ (Theorem˜3.16). We also obtain positivity of the coefficients of these univariate restrictions for strictly $\mathcal{K}$}-CLC polynomials (Proposition˜3.14, Corollary˜3.15).
Then we connect these structural inequalities to classical Hurwitz stability of univariate polynomials. Using Routh–Hurwitz and Liénard–Chipart criteria together with the log-concavity relations among coefficients, we show that strict $\mathcal{K}$}-CLC implies Hurwitz stability in low degrees. Specifically, we prove that every univariate strictly CLC polynomial of degree $d\leq 4$ with positive coefficients is Hurwitz-stable (Proposition˜4.14), and we obtain refined sufficient conditions for Hurwitz stability in degree $d=5$ (Theorem˜4.17) based on strengthened log-concavity inequalities.
In the final part of this paper our Newton-type coefficient inequalities for strictly CLC and strictly $\mathcal{K}$}-Lorentzian polynomials yield simple sufficient tests ensuring stability of a linear time-invariant system (LTI) directly from the coefficients of $\chi_{A}(t)$, (Theorem˜5.2 and Theorem˜5.4).
Organization. The rest of this paper is structured as follows. In Section˜2 we recall basic definitions on proper cones, $\mathcal{K}$}-nonnegativity, $\mathcal{K}$}-irreducibility, and the equivalence between $\mathcal{K}$}-Lorentzian and $\mathcal{K}$}-CLC forms. In Section˜3 we develop several key properties of $\mathcal{K}$}-Lorentzian polynomials, including Hessian characterizations, cone-theoretic Perron–Frobenius conditions, and a line-restriction framework for $\mathcal{K}$}-CLC, leading to log-concavity of directional derivatives. In Section˜4 we relate these inequalities to the classical half-plane property and Hurwitz stability for univariate restrictions, deriving degree-dependent criteria for stability. Section˜5 applies the Hurwitz–stability criteria and Newton-type log-concavity inequalities developed earlier to the characteristic polynomials of linear time-invariant (LTI) systems, yielding simple sufficient conditions for (asymptotic) stability expressed directly in terms of their coefficients.
@@ 2. State of the Art
Let ${\mathbb{R}}[x]$ represent the space of $n$-variate polynomials over $\mathbb{R}$}, and ${\mathbb{R}}[x]_{\leq d}$ represent the space of $n$-variate polynomials over $\mathbb{R}$} with degree at most $d$ and ${\mathbb{R}}[x]^{d}_{n}$ denote the set of real homogeneous polynomials (aka forms) in $n$ variables of degree $d$. A polynomial $f\in{\mathbb{R}}[x]$ is said to be (strictly) log-concave at a point $a\in{\mathbb{R}}^{n}$ if $f(a)>0$, and $\log f$ is a (strictly) concave function at $a$, i.e. the Hessian of $\log f$ is negative semidefinite (negative definite) at $a$. A polynomial $f\in{\mathbb{R}}[x]$ is log-concave on a proper cone ${\mathcal{K}}\subset{\mathbb{R}}^{n}$ if $f$ is log-concave at every point of the interior of $\mathcal{K}$}. By convention, the zero polynomial is log-concave (but not strictly log-concave) at all points of ${\mathbb{R}}^{n}$.
If ${\mathcal{K}}={\mathcal{K}}^{\ast}$, a self-dual cone, then $K$ is closed, pointed, and full dimensional[Barkerselfdual]. We denote quadratic forms by a lower case letter, and the matrix of the quadratic form by the corresponding upper case letter, i.e. if $q$ is a quadratic form, then its matrix is $Q$ and $q(x)=x^{t}Qx$.
@@
[GPlorentzian] A polynomial (form) $f\in{\mathbb{R}}[x]_{\leq d}$ is called a $\mathcal{K}$}-completely log-concave aka $\mathcal{K}$}-CLC (form) on a proper convex cone $\mathcal{K}$} if for any choice of $a_{1},\dots,a_{m}\in{\mathcal{K}}$, with $m\leq d$, we have that $D_{a_{1}}\dots D_{a_{m}}f$ is log-concave on $\operatorname{int}{\mathcal{K}}$. A polynomial (form) $f\in{\mathbb{R}}[x]$ is strictly $\mathcal{K}$}-CLC if for any choice of $a_{1},\dots,a_{m}\in{\mathcal{K}}$, with $m\leq d$, $D_{a_{1}}\dots D_{a_{m}}f$ is strictly log-concave on all points of $\mathcal{K}$}.
@@
[GPlorentzian] Let $\mathcal{K}$} be a proper convex cone. A form $f\in{\mathbb{R}}[x]_{n}^{d}$ of degree $d\geq 2$ is said to be $\mathcal{K}$}-Lorentzian if for any $a_{1},\dots,a_{d-2}\in\operatorname{int}{\mathcal{K}}$, the quadratic form $q=D_{a_{1}}\dots D_{a_{d-2}}f$ satisfies the following conditions:
@@
Let $\mathcal{K}$} be a proper convex cone in ${\mathbb{R}}^{n}$. An $n\times n$ matrix $A$ is $\mathcal{K}$}-nonnegative if $Ax\in{\mathcal{K}}$ for any $x\in{\mathcal{K}}$, i.e. $A({\mathcal{K}})\subseteq{\mathcal{K}}$. A matrix $A$ is $\mathcal{K}$}-positive if $Ax\in\operatorname{int}{\mathcal{K}}$ for any nonzero $x\in{\mathcal{K}}$. A $\mathcal{K}$}-nonnegative matrix $A$ is called $\mathcal{K}$}-irreducible if $A$ leaves no (proper) face of $\mathcal{K}$} invariant. A $\mathcal{K}$}-nonnegative matrix which is not $\mathcal{K}$}-irreducible is called $\mathcal{K}$}-reducible, cf. [gantmacher], [Vandergraftspectral].
@@
[GPlorentzian, Lemma 3.1] A quadratic form $q(x)=x^{t}Qx\in{\mathbb{R}}[x]_{n}^{2}$ is $\mathcal{K}$}-Lorentzian if and only if the matrix $Q$ of $q$ has exactly one positive eigenvalue and one of the following conditions holds:
@@
Similar to copositive matrices over the nonnegative orthant, we define $\mathcal{K}$}-copositive matrices over $\mathcal{K}$}, see [stabilitylevi]. A matrix $Q$ is (strictly) $\mathcal{K}$}-copositive if $x^{t}Qx(>0)\geq 0$ for all $x\in{\mathcal{K}}$.
@@
[GPlorentzian, Cor 3.8] Let $\mathcal{K}$} be a self-dual cone. Then the quadratic form $q=x^{t}Qx$ is $\mathcal{K}$}-Lorentzian if and only if the matrix $Q$ of $q$ has exactly one positive eigenvalue and $Q$ is $\mathcal{K}$}-nonnegative.
@@
[GPlorentzian, Cor 3:10] Let ${\mathcal{K}}\subset{\mathbb{R}}^{n}$ be a self-dual cone. Then $q=x^{t}Qx$ is $\mathcal{K}$}-Lorentzian if and only if the matrix $Q$ of $q$ has exactly one positive eigenvalue and $Q$ is either nonsingular and $\mathcal{K}$}-irreducible, or singular and $\mathcal{K}$}-nonnegative.
@@
Furthermore, we use the following equivalence result between $\mathcal{K}$}-Lorentzian and $\mathcal{K}$}-CLC polynomials, proved in [GPlorentzian, Theorem 4:10].
Theorem 2.9.
Let ${\mathcal{K}}\subset{\mathbb{R}}^{n}$ be a proper cone. A form $f\in{\mathbb{R}}[x]_{n}^{d}$ is $\mathcal{K}$}-Lorentzian if and only if $f$ is $\mathcal{K}$}-CLC. Equivalently, $\operatorname{\mathrm{SL}}^{d}({\mathcal{K}})=\mathrm{CLC}^{d}({\mathcal{K}})$.